# GEIC: Universal and Multilingual Named Entity Recognition with Large Language Models

### v2

@@ 5.2 Results
The results indicate that when handling NER tasks with ample training resources and simple classifications, LLM-based methods still lag behind existing methods, whether on the English-only CoNLL2003 or the multilingual PAN-X. However, in low-resource scenarios, LLM-based methods, whether supervised or few-shot, achieve significantly better results. CascadeNER surpasses existing methods on CrossNER except Music, and FewNERD, highlighting its exceptional generalization and capability to handle complex entity categorization.
However, some challenges remains. Although the GEIC task is designed to be able to accommodate nested and discontinuous NER, CascadeNER is currently limited to flat NER tasks. This limitation arises from the fact that the models in CascadeNER are pre-trained on the GEIC version of AnythingNER, and there is a significant lack of suitable multilingual corpora for nested and discontinuous NER. Our resources are insufficient to collect enough open-source data for this purpose, which lead to AnythingNER containing only flat NER labels, and thus constraining CascadeNER to flat NER. Furthermore, due to resource limitations, AnythingNER currently supports only 8 languages, excluding other cucial languages, somewhat restricting its applicability.
6 Conclusion
This paper introduces Generation-based Extraction and In-context Classification (GEIC), a novel task which is designed for applying LLMs in NER tasks, especially few-shot and zero-shot scenerios. The idea of GEIC can benefit a wide range of scenarios, not only the LLM-based NER tasks, but also inspiring researchers to apply relatively small LLMs in other complex NLP tasks, such as Relation Extraction. To demonstrate GEIC, we propose CascadeNER, the first GEIC framework, and AnythingNER, the first dataset for LLM-based NER tasks like GEIC. Extensive experiments show that CascadeNER excels in low-resource and complex classification scenarios, demonstrating strong practicality. We hope that our work will inspire further research on the application of LLMs in NLP tasks.
Appendix A Dynamic Categorization
In addition to the BIO and GEIC format with the traditional fixed categorization for AnythingNER, we introduce a GEIC version using dynamic categorization. Dynamic categorization is a novel categorization method designed for LLM-based NER training and evaluation, particularly suited for GEIC. In the dynamic categorized version of AnythingNER, we dynamically adjust the labels and corresponding entity type list for entities during annotation, overcoming the constraints of static classification. By flexibly altering the granularity and numbers of categories during data annotation, this method effectively trains and evaluates the model’s generalization. Adjustments include mixing types of different granularities, using type lists without irrelevant types, replaced types with synonyms, and and merging certain types into miscellaneous/others, as shown in Figure 6. This approach revolutionarily addresses the issue of traditional datasets being unsuitable for few-shot and zero-shot training and evaluation, simulating real-world scenarios. CascadeNER, trained on this version of the dataset, exhibits significantly improved generalization ability with extraordinary few-shot and zero-shot performance.
Figure 6: Examples of dynamic categorization methods. The red box contains the fine-grained categories of the entity in standard categorization. The green boxes contains the category lists modified with four dynamic categorization methods. The blue boxes represents the entity category corresponding to each category list.
To mitigate potential consistency issues arising from dynamic categorization, we provided annotators with detailed guidelines. Following annotation, we conduct statistical analysis and adjust the dynamically categorized entities to maintain a balance between generalization and consistency. Notably, as dynamic categorization is a subtractive approach applied to a comprehensive classification system, the success of this method relys heavily on AnythingNER’s comprehensive categorization system, which includes 155 entity types. This approach may not be suitable for all datasets.
Appendix B Detail of AnythingNER
In this section, we briefly present some key metrics of AnythingNER and compare AnythingNER with existing datasets. Table 4 presents a simple comparison between AnythingNER and existing multilingual or fine-grained datasets.
Dataset #Language #Coarse #Fine Domain
CoNLL2002 2 4 no News
CoNLL2003 2 4 no News
ACE2005 2 7 41 News
OntoNotes 5.0 3 18 no General
WNUT2017 1 6 no SocialMedia
FEW-NERD 1 8 66 General
PAN-X 282 3 no General
MultiCoNER 11 6 33 General
I2B2 1 22 no Medical
AnythingNER 8 8 155 General
Table 4: Overview of NER datasets. Notably, AnythingNER covers a wide range of cross-domain categories, such as art, medicine, and biology, thus offering better generalization compared to other general datasets.
The specific data volumes for each language are shown in Table 5. It is important to note that for languages except English and Chinese, we partially use manually translated English corpora. This is necessary to balance category distribution, as most languages lack sufficient corpora in specific domains. We also provide conversion scripts that allow AnythingNER to be transformed into train, dev, and test sets with non-overlapping subsets based on coarse categories, making it easier to use for few-shot methods.
Language # Sentences # Tokens # Entities #Train #Dev #Test
English 1500 27.3k 3709 900 300 300
Chinese 1500 26.4k 3675 900 300 300
Spanish 1000 18.6k 2393 602 201 197
French 1000 19.2k 2475 600 200 200
German 1000 18.8k 2473 603 197 200
Japanese 1000 18.4k 2387 600 199 201
Korean 1000 17.9k 2376 598 200 202
Russian 1000 19.3k 2406 601 198 201
Table 5: Statistics of AnythingNER across languages. We roughly follow a 3:1:1 ratio to divide the train, dev, and test sets, with slight adjustments based on the proportional distribution of entities within the corpus.
Appendix C Ablation Study
C.1 Result Fusion
@@
In Section 3.2.1, we introduce our union strategy in result fusion to address the issue of extractor recall being significantly lower than precision, allowing multiple extractions for one sentence and taking the union of the results to maximize recall. For the problem of entity nesting, where different extraction rounds yield overlapping or nested entities, we adopt a length-first strategy, retaining the longer entity. Table 6 provides a example for the significantly low recall.
@@
Table 6: Precision, recall, and F1 Score for CoNLL2003 and CrossNER. In this experiment, both base models used in CascadeNER are Qwen2-1.5B, and the results are obtained in zero-shot scenarios.
Figure 7 presents the impact of increasing the number of extraction repetitions in zero-shot scenarios on CoNLL2003. The results show that our strategy can slightly improve recall with minimal impact on precision. Given the obvious margin effect after 3 repetitions, we ultimately select $3$ as the repetition count $k$ for other experiments. It is important to emphasize that even without repeated extraction, CascadeNER demonstrates a significant performance advantage.
Figure 7: The curves showing visualized precision, recall, and F1 Score as a function of the number of repetitions, demonstrating how these metrics change with increasing repetition counts $k$. In this experiment, both base models used in CascadeNER are Qwen2-1.5B.
C.2 Varying the Base Models
In this section, we use four different SLMs as the base models for CascadeNER, i.e., two versions of Qwen2 and Gemma with different parameters, namely Qwen2-1.5B, Qwen2-7B, Gemma-2B, and Gemma-7B. These models represent the current best-performing SLMs. We use the dynamic categorized version of AnythingNER to fine-tune the instruct versions of the four models. Each model is fine-tuned separately on the corresponding dataset to obtain both an extractor and a classifier. In addition to using combinations where the extractor and classifier are based on the same original model, we also set up combinations where the extractor uses the 1.5B/2B model and the classifier uses the 7B model, resulting in a total of six combinations. The performance comparison of these combinations on AnythingNER is shown in Table 7. Based on these results, Qwen2 outperform Gemma in multilingual tasks overall, and the performance of Qwen2-1.5B is slightly higher than Gemma-2B, while significantly reducing resource consumption. Therefore, we chose Qwen2 as the base models for other experiments.
English Chinese Spanish Japanese Macro F1
Qwen2-1.5B 71.2 67.5 65.3 65.7 67.4
Qwen2-combine 73.2 70.4 67.2 66.2 69.2
Qwen2-7B 73.3 71.8 68.1 67.7 70.2
Gemma-2B 72.9 61.5 69.1 62.4 66.5
Gemma-combine 74.6 63.3 69.7 64.2 68.0
Gemma-7B 74.8 65.9 68.3 65.1 68.5
Table 7: Comparison of F1 scores for CascadeNER using different models on AnythingNER. The results are obtained in zero-shot scenarios.
C.3 Context in Classification
@@
Table 8: Both base models used in CascadeNER are Qwen2-1.5B. The results are obtained in zero-shot scenarios. The used datasets are CoNLL2003 and CrossNER. Accuracy is used in the evaluation of classifiers.
C.4 Number of Few-shot Demonstrations
In this section, we conduct experiments to estimate the effect of the number of demonstrations. We compare CascadeNER’s performance in the few-shot scenario on the CoNLL2003 dataset as the number of demonstrations increases with the current SOTA method for this dataset, ACE+document-context. The results are shown in Figure 9. While CascadeNER already demonstrate excellent performance in the zero-shot scenario, its performance improves further when provided with a certain number of few-shot demonstrations. CascadeNER shows a significant advantage in low-resource scenarios, but as the number of demonstrations increases, this advantage gradually diminishes until it is eventually surpassed by a tiny difference.
Figure 9: Comparison between CascadeNER and ACE+document-context. Both base models used in CascadeNER are Qwen2-1.5B. Demonstrations, or the sentences in the train set, are start from 0 to 10000. The sampling follows the algorithm in Section 3.2.3.
Appendix D Comparsion of Prompt
@@
PromptNER utilizes detailed descriptions of each entity’s specific definition and reasoning processes to fully leverage the LLM’s logical reasoning abilities. However, like traditional methods, it treats NER as a sequence labeling task, failing to effectively utilize the LLM’s global contextual understanding capabilities, making it prone to overlooking important context in complex sentences. Additionally, the task descriptions are overly complex, which not only makes it difficult for the model to correctly execute tasks in zero-shot scenarios, but also leads to a higher likelihood of hallucinations in tasks requiring fine-grained classification, such as the fine-grained settings of FewNERD and AnythingNER, as each category’s definition requires detailed descriptions. These issues reduce PromptNER’s generalization and accuracy, limiting its application.
@@
CascadeNER is based on the concept of the GEIC task, dividing the NER task into two sub-tasks: extraction and classification, while simplifying the input and output formats and reducing logical complexity. This ensures that even SLMs with limited capacity to handle complex tasks can still perform the tasks accurately and efficiently. In extraction, CascadeNER leverages the model’s generation capabilities by producing sentences with identifiers, treating all entities uniformly, which enhances the model’s generalization ability across different languages and domains. Notably, it avoids reliance on word order by consistently using "##" to mark entities, ensuring consistent annotation regardless of whether the language is right-to-left or left-to-right, improving cross-language consistency and adaptability. In classification, our method processes the entire sentence as a whole, better utilizing LLMs’ strengths in contextual understanding and semantic modeling. By leveraging the LLM’s ability to model long-range dependencies, the model’s capacity to handle complex sentence structures is enhanced, avoiding fragmentation of information and improving overall consistency and generalization. However, our method also has limitations. The use of unified identifiers prevents CascadeNER from effectively handling nested NER. We plan to address this by developing a solution that accommodates both multilingual and nested NER tasks in future.
Appendix E Data Contamination Statement
@@
Appendix F Ethical Statement of AnythingNER
When constructing AnythingNER, we strictly adhere to existing ethical guidelines Bender and Friedman (2018); Gebru et al. (2021); Hovy and Spruit (2016), ensuring that our data sources and processing methods comply with legal and ethical standards while maintaining high-quality annotations. All the text in AnythingNER is sourced from Wikipedia, ensuring no violations of privacy or copyright, as Wikipedia is an open-source platform with user-contributed content from around the world. During data collection and annotation, we balance category distribution to minimize the risk of bias in the model. Furthermore, we maintain transparency by detailing the dataset development process and data partitioning in this paper, ensuring clarity and reproducibility for future research.
For the annotators, each language in AnythingNER is annotated by two junior or higher-level students from the corresponding language departments at our university. Due to the double-blind review process, the annotators’ identities cannot be disclosed in this version. Each annotator receives extensive training and follows AnythingNER’s multi-granularity classification system to ensure consistent and accurate entity annotations across various languages and domains. The annotation process for each language are divided into two parts equally, with each annotator independently handling one part. After the initial annotation, GPT-4o OpenAI (2023) is utilized for machine review, and the annotators revise their work based on the review results. For ambiguous terms or specialized domain terms, the annotators either collaborate with each other or consult experts to ensure the accuracy and reliability of the annotations.
Appendix G Detailed Categories of AnythingNER
G.1 Person
@@ G.8 Science Entity
Appendix H Best Results of CascadeNER
CascadeNER 93.5 76.5 76.6 83.9 83.6 78.2 86.3 76.3
Table 9: Best F1 score of CascadeNER on CoNLL2003, CrossNER, and FewNERD.
Model PAN-X MultiCoNER
en es fr ru de zh ja ko en es ru de zh ko
CascadeNER 92.1 86.1 87.5 87.1 83.8 87.9 84.4 81.4 87.1 82.2 80.7 70.5 86.2 78.3
Table 10: Best F1 score of CascadeNER on PAN-X and MultiCoNER.

### v4

@@ D.4 Experimental Results
As shown in Table 6 and 7, the results indicate that in low-resource scenarios, LLM-based methods achieve significantly better results. CascadeNER surpasses existing methods on CrossNER except Music and FewNERD, and PAN-X and MultiCoNER in some languages, achieving new SOTA performance and highlighting its exceptional generalization and capability to handle complex entity categorization. However, when handling NER tasks with ample training resources and simple classifications, LLM-based methods still lag behind existing methods, whether on the English-only CoNLL2003 or the multilingual PAN-X, indicating that supervised methods are still useful in some scenerios.
Appendix E Ablation Study
E.1 Result Fusion
@@
In Section 4.2, we introduce our union strategy in result fusion to address the issue of extractor recall being significantly lower than precision, allowing multiple extractions for one sentence and taking the union of the results to maximize recall. For the problem of entity nesting, where different extraction rounds yield overlapping or nested entities, we adopt a length-first strategy, retaining the longer entity. Table 8 provides a example for the significantly low recall.
@@
Table 8: Precision, recall, and F1 Score for CoNLL2003 and CrossNER. In this experiment, both base models used in CascadeNER are Qwen2.5-7B, and the results are obtained in zero-shot scenarios.
Figure 7 presents the impact of increasing the number of extraction repetitions in zero-shot scenarios on CoNLL2003. The results show that our strategy can slightly improve recall with minimal impact on precision. Given the obvious margin effect after 3 repetitions, we ultimately select $3$ as the repetition count $k$ for other experiments. It is important to emphasize that even without repetition, CascadeNER still has a significant performance advantage.
Figure 7: The curves showing visualized precision, recall, and F1 Score as a function of the number of repetitions, demonstrating how these metrics change with increasing repetition counts $k$. Both base models used in CascadeNER are Qwen2.5-7B.
E.2 Varying the Base Models
In this section, we use four different lightweight LLMs as the base models for CascadeNER, i.e., two versions of Qwen2 and Gemma with different parameters, namely Qwen2-1.5B, Qwen2-7B, Gemma-2B, and Gemma-7B. Gemma is another prominent lightweight LLM series, proposed by Google Team et al. (2024). These models represent the current best-performing lightweight LLMs. We use the dynamic categorized version of DynamicNER to fine-tune the instruct versions of the four models. Each model is fine-tuned separately on the corresponding dataset to obtain both an extractor and a classifier. The performance comparison of these combinations on a selected part of DynamicNER is shown in Table 9. Based on these results, Qwen2.5 outperform Gemma in multilingual tasks overall. Therefore, we choose Qwen2.5 as the base models for other experiments.
English Chinese Spanish Japanese
Qwen2.5-1.5B 62.8 58.9 55.7 54.1
Qwen2.5-7B 68.2 64.5 61.5 60.8
Gemma-2B 58.9 49.5 53.1 48.4
Gemma-7B 61.7 53.9 55.3 52.1
Table 9: F1 scores for CascadeNER with different base models on a selected part of DynamicNER
E.3 Context in Classification
@@
Table 10: Both base models used are Qwen2.5-7B. The results are obtained in zero-shot scenarios. The used datasets are CoNLL2003 and CrossNER. Accuracy is used in the evaluation of classifiers.
Appendix F LLM-based Methods Comparsion
@@
PromptNER utilizes detailed descriptions of each entity’s specific definition and CoT reasoning processes to fully leverage the LLM’s logical reasoning abilities. However, like traditional methods, it treats NER as a sequence labeling task, failing to effectively utilize the LLM’s global contextual understanding capabilities, making it prone to overlooking important context in complex sentences. Additionally, the task descriptions are overly complex, which not only makes it difficult for lightweight LLMs to correctly execute tasks, but also leads to a higher likelihood of hallucinations in tasks requiring fine-grained classification, such as the fine-grained settings of FewNERD and DynamicNER, as each category’s definition requires descriptions. These issues reduce PromptNER’s generalization and accuracy, limiting its application.
@@
CascadeNER divides the NER task into two sub-tasks: extraction and classification, while simplifying the input and output formats and reducing logical complexity. This ensures that even lightweight LLMs with limited capacity to handle complex tasks can still perform the tasks accurately and efficiently. In extraction, CascadeNER leverages the model’s generation capabilities by producing sentences with identifiers, treating all entities uniformly, which enhances the model’s generalization ability across different languages and domains. Notably, it avoids reliance on word order by consistently using "##" to mark entities, ensuring consistent annotation regardless of whether the language is right-to-left or left-to-right, improving cross-language consistency and adaptability. In classification, our method processes the entire sentence as a whole, better utilizing LLMs’ strengths in contextual understanding and semantic modeling. By leveraging the LLM’s ability to model long-range dependencies, the model’s capacity to handle complex sentence structures is enhanced, avoiding fragmentation of information and improving overall consistency and generalization. However, our method also has limitations. The use of unified identifiers prevents CascadeNER from effectively handling nested NER. We plan to address this by developing a solution that accommodates both multilingual and nested NER tasks in future.
Appendix G Computational Resource Usage Record
In Table 11, we provide the API costs incurred when testing the complete dynamic version of DynamicNER in few-shot scenarios using three LLM-based methods with GPT-4o, serving as a reference for the computational resources required by these methods. The cost calculation follows OpenAI’s official GPT-4o pricing, with input costs at 2.5 USD per 1M tokens and output costs at 10 USD per 1M tokens. The records show that CascadeNER exhibits significant advantages over existing methods in computational resource consumption.
GPT-NER PromptNER CascadeNER
Cost (USD) 513.92 128.49 45.86
Table 11: Cost comparison of three LLM-based methods. The cost is calculated according to OpenAI’s official GPT-4o pricing, not the actual cost.
Appendix H Data Contamination Statement
@@
In additional experiments on CascadeNER, we notice another critical data contamination concern: potential corpus overlap between DynamicNER and other benchmark datasets utilizing Wikipedia-derived text, which can reduce evaluation fairness. To mitigate this risk, we implement a rigorous filtering protocol during DynamicNER’s annotation phase. After completing the initial manual annotation of the base version, we employ Sentence-BERT to compute semantic cosine similarity between each candidate sentence and existing sentences in reference datasets. Sentences exhibiting similarity scores exceeding 0.8 are excluded from the corpus. New sentences from collected corpus meeting the similarity criteria are then re-annotated following the original annotation workflow. This iterative process continues until all sentences in the base version satisfy the similarity constraints. After this we utilize dynamic categorization to generates the dynamic version. This procedure ensures the reliability and fairness of our test results.
Appendix I Ethical Statement of DynamicNER
When constructing DynamicNER, we strictly adhere to existing ethical guidelines Bender and Friedman (2018); Gebru et al. (2021); Hovy and Spruit (2016), ensuring that our data sources and processing methods comply with legal and ethical standards while maintaining high-quality annotations. All the text in DynamicNER is sourced from Wikipedia, ensuring no violations of privacy or copyright, as Wikipedia is an open-source platform with user-contributed content from around the world. During data collection and annotation, we balance category distribution to minimize the risk of bias in the model. Furthermore, we maintain transparency by detailing the dataset development process and data partitioning in this paper, ensuring clarity and reproducibility for future research.
For the annotators, each language in DynamicNER is annotated by two junior or higher-level students from the corresponding language departments at our university. Due to the double-blind review process, the annotators’ identities cannot be disclosed in this version. Each annotator receives extensive training and follows DynamicNER’s multi-granularity classification system to ensure consistent and accurate entity annotations across various languages and domains. The annotation process for each language are divided into two parts equally, with each annotator independently handling one part. After the initial annotation, the annotators revise their work based on the review results. For ambiguous terms or specialized domain terms, the annotators either collaborate with each other or consult experts to ensure the accuracy and reliability of the annotations.
In our writing, we use CahtGPT-4o Achiam et al. (2023) and Claude 3.5 Anthropic (2023) for assistance.
Appendix J Detailed Categories of DynamicNER
J.1 Person
@@ Appendix I Ethical Statement of DynamicNER
Appendix K More Categorization Quality Evaluation
In this section, we display the quantitative results of categorization metrics in Spanish, French, Russian, German, Japanese, and Korean. The results in shown in Figure 10. Experimental results demonstrate that our dynamic categorization method maintains or improves dataset quality compared to the base version in all languages.
Figure 10: Quantitative categorization metric results for 3 versions DynamicNER.